# Deformable Endoscopic Tissues Reconstruction with Gaussian Splatting

### v1

@@ 3.2 Implementation Details
In our approach, we adopt the two-stage training methodology in [26] to model the static and deformation fields. In the first stage, we train 3D Gaussian models for the static field, while in the second stage, we focus on training the deformation field. We conduct 3,000 and 60,000 iterations for the first and second stages, respectively. The initial point clouds are estimated using COLMAP [21]. The importance maps scaling strength $\alpha$ is set to 30 and $\delta$ = 0.2 for depth loss. Hyperparameters $\lambda_{D}$, $\lambda_{TV1}$, and $\lambda_{TV2}$ are set to 0.03, 0.01, and 1.0.
@@ 3.3 Results
Table 1: Quantitative comparison on rendering quality and speed.
@@
EndoNeRF [25] 35.624 0.942 0.064 $<0.2$
ForPlane [27] 36.457 0.946 0.058 $\sim 1.7$
EndoGS (Ours) 37.654 0.965 0.036 $\sim 40$
@@
We compare EndoGS against two methods, i.e., EndorNerf [25] and ForPlane [27] (an updated version of LerPlane [28]), due to their competitive rendering quality. ForPlane is trained for 32k iterations. We use the same masks in training and evaluation for three methods and evaluate the rendering results on the same cropped zone where the lowest part of the videos that contain display patterns are removed.
Figure 3: Ablation on depth regularization. We show rendering depth maps with different depth regularization strength on scene "pulling soft tissues". The values in the bracket denote the weight of the depth regularization loss.
@@
Fig. 2 presents a qualitative comparison on scene "traction" between EndoGS and competitive baselines. On the rendering quality, EndoGS clearly outperforms other methods. Tissues deformation occurs at different timesteps, and our method better reconstructs the deformation along the time. Table 1 presents a quantitative comparison on the 6 videos, showcasing the superior performances of our method over the baseline methods in terms of various evaluation metrics related to rendering quality and speed. Considering that EndoNeRF exhibits poor performance on the first rendering frame, we exclude the first frames from the evaluation process for all three methods. In terms of rendering quality, EndoGS demonstrates a significant advantage, outperforming the other methods by a considerable margin. Furthermore, EndoGS benefits from the rendering efficiency of Gaussian Splatting, enabling it to achieve real-time rendering speeds. In contrast, the baseline methods struggle to maintain a high FPS rate, highlighting the superiority of EndoGS in this regard.
Ablation Analysis. Fig. 3 presents the ablation study on depth regularization. While we optimize the Gaussians from a single-viewpoint video, 3D information is missing and it is likely the representations are overfitted. We leverage estimated depth maps to help mitigate the illness. It is shown that without depth regularization or with a smaller weight of the regularization, the rendered depth maps are strayed away from the depth guidance. In Fig. 4, we show the effectiveness of spatial Total Variation loss. While we optimize the Gaussians on seen pixels in frames with the tool masks, the lack of continuity leads to color drift in rendering results (frame w/o spatial TV). This phenomenon can be mitigated during the optimization with spatial TV loss.
@@ 4 Conclusion
We present a method based on Gaussian Splatting for deformable endoscopic tissues reconstruction, rendering high-quality deformable tissues in real-time from a single-viewpoint video, estimated depth maps, and labeled tool masks. Experiments on DaVinci robotic surgery videos show higher rendering quality.
Limitations and Future Works. Our work has limitations in two folds. First, 3D reconstruction from single-viewpoints videos is an ill-posed problem, making it infeasible for surgical downstream applications. Second, due to the lack of 3D cues and surface constrains in single-viewpoint videos for Gaussian Splatting optimization, it is likely that artifacts and ambiguities happen in 3D reconstruction with novel viewpoints. Accordingly, future works can focus on practical endoscopic reconstruction with surface-aligned Gaussian Splatting. 3D tissues reconstruction with more surgical cameras is needed to facilitate realistic downstream tasks. Additionally, regularization that encourages the Gaussians to well distributed over the surface can mitigate weaknesses on reconstructed scene surface and geometry, as suggested by [7].
@@ Acknowledgement
We thank Med-AIR Lab CUHK for DaVinci robotic prostatectomy data.

### v3

@@ 3.2 Implementation Details
In our approach, we adopt the two-stage training methodology in [30] to model the static and deformation fields. In the first stage, we train 3D Gaussian models for the static field, while in the second stage, we focus on training the deformation field. We conduct 3,000 and 60,000 iterations for the first and second stages, respectively. The initial point clouds are estimated using COLMAP [25]. The importance maps scaling strength $\alpha$ is set to 30 and $\delta$ = 0.2 for depth loss. Hyperparameters $\lambda_{D}$, $\lambda_{TV1}$, $\lambda_{TV2}$, $\lambda_{S}$ are set to 0.5, 0.01, 1.0, 0.2. We train our models on an NVIDIA RTX 3090 GPU.
@@ 3.3 Results
Table 1: Quantitative comparison on rendering quality and speed (standard deviation in parentheses).
@@
EndoNeRF [29] 35.112 (1.470) 0.936 (0.021) 0.066 (0.030) $<0.2$
ForPlane [32] 36.427 (1.214) 0.947 (0.007) 0.058 (0.003) $\sim 1.7$
EndoGS w/o SA 37.603 (0.076) 0.964 (0.005) 0.036 (0.009) $\sim 70$
EndoGS (Ours) 37.935(0.088) 0.966(0.003) 0.034(0.001) $\sim 70$
@@
We compare EndoGS against two methods, i.e., EndoNeRF [29] and ForPlane [32] (an updated version of LerPlane [33]), due to their competitive quality. ForPlane is trained for 32k iterations. We use the same masks in training and evaluation for three methods and evaluate the rendering results on the same cropped zone where the lowest part of the videos that contain display patterns are removed.
@@
Figure 4: Ablation on Surface-Aligned regularization. We show reconstruction results w/o and w/ Surface-Aligned regularization on scene "pulling soft tissues".
Fig. 2 presents a qualitative comparison on scene "traction" between EndoGS and competitive baselines. On the rendering quality, EndoGS clearly outperforms other methods. Tissues deformation occurs at different timesteps, and our method better reconstructs the deformation along the time. We also include the ablation results over the Surface-Aligned regularization which is denoted as SA in the table. Table 1 presents a quantitative comparison on the 6 videos, showcasing the superior performances of our method over the baseline methods in terms of various evaluation metrics related to rendering quality and speed. Considering that EndoNeRF exhibits poor performance on the first rendering frame, we also exclude the first frames from the evaluation process. In terms of rendering quality, EndoGS demonstrates a significant advantage, outperforming the other methods by a considerable margin. Furthermore, EndoGS benefits from the rendering efficiency of Gaussian Splatting, enabling it to achieve real-time rendering speeds. In contrast, the baseline methods struggle to maintain a high FPS rate, highlighting the superiority of EndoGS in this regard.
Ablation Analysis. In Fig. 3, we show the effectiveness of spatial Total Variation loss. While we optimize the Gaussians on seen pixels in frames with the tool masks, the lack of continuity leads to color drift in rendering results (frame w/o spatial TV). This phenomenon can be mitigated during the optimization with spatial TV loss. Table 1 and Fig. 4 present the quantitative and qualitative results over the effectiveness of Surface-Aligned regularization. When reconstructing the surrounding of tool occlusion masks, noticeable artifacts on the surface are evident, which cause the rendered image to become blurred. It is shown that with Surface-Aligned regularization more details are preserved.
@@ 4 Conclusion
We present a method for deformable endoscopic tissue reconstruction based on surface-aligned Gaussian splatting, which achieves high-quality, real-time reconstruction from a single-viewpoint video, estimated depth maps, and labeled tool masks. Experiments on DaVinci robotic surgery videos demonstrate the superior quality of our method. However, there is limitation to 3D reconstruction from single-viewpoint videos, as it is an ill-posed problem that makes it infeasible for surgical downstream applications. For example, we lack 3D cues for tissues that are occluded by tools. Therefore, future work can focus on practical endoscopic reconstruction, and 3D tissue reconstruction with more surgical cameras is needed to facilitate realistic downstream tasks. We would also like to recommend readers referring to other concurrent and follow-up works [16, 9, 1, 36, 38, 34, 11, 15].
@@ Acknowledgement
This work was partially supported by the Research Grants Council of Hong Kong ( T45-401/22-N and 27206123) and the National Natural Science Foundation of China (No. 62201483). We thank Med-AIR Lab CUHK for DaVinci robotic prostatectomy data.